# Identification and characterization of distorted gravitational waves by lensing using deep learning

### v1

@@ Abstract
Gravitational waves (GWs) can be distorted by intervening mass distributions while propagating, leading to frequency-dependent modulations that imprint a distinct signature on the observed waveforms. Bayesian inference for GW lensing with conventional sampling methods is costly, and the problem is exacerbated by the rapidly growing GW catalog. Moreover, assessing the statistical significance of lensed candidates requires thousands, if not millions, of simulations to estimate the background from noise fluctuations and waveform systematics, which is infeasible with standard samplers. We present a novel method, DINGO-lensing, for performing inference on lensed GWs, extending the neural posterior estimation framework DINGO. By comparing our results with those using conventional samplers, we show that the compute time of parameter estimation of lensed GWs can be reduced from weeks to seconds, while preserving accuracy both in the posterior distributions and the evidence ratios. We train our neural networks with LIGO detector noise at design sensitivity and a lens model that accommodates two overlapping chirps with opposite parity. We show that the lensing parameters are recovered with millisecond precision for the time delays. We also demonstrate that our network can identify signals diffracted by point masses, highlighting its flexibility for searches. By simulating thousands of lensed and nonlensed events, we determine how the detectability changes with different source properties. DINGO-lensing provides a scalable and efficient avenue for identifying and characterizing gravitationally lensed GW events in the upcoming observing runs.
@@ I Introduction
GWs propagate over cosmological distances and are not impervious to intervening matter. Instead, their path is affected by nearby gravitational potentials (acting as lenses), such as those of stars, black holes, galaxies, and clusters of galaxies, leading to gravitational lensing Nakamura (1998); Takahashi and Nakamura (2003). In strong gravitational lensing, multiple signals, or images, of the source are produced each with a distinct amplitude, phase shift, and time delay Schneider et al. (1992). The arrival time difference between these images can be very small for smaller lens masses (or specific source-lens-observer configurations) Schneider et al. (1992); Lo et al. (2025); Vujeva et al. (2025a); Ezquiaga et al. (2025); Vujeva et al. (2025b), making it challenging to look for signals coming from the same source. Additionally, these small lens masses may produce overlapping images which causes a beating pattern in the inspiral portion of the waveform resembling precession or the effect of two different binary signals overlapping Liu and Kim (2024); Rao et al. (2025). In the full wave-optics regime, it is also possible to have a single signal with frequency-dependent modulations due to diffraction effects. Detecting such signal can be challenging in matched-filtering searches Chan et al. (2025a), but it can answer key questions in the small lens mass regime regarding small compact lens populations Jung and Shin (2019); Lai et al. (2018); Christian et al. (2018); Dai et al. (2018); Diego et al. (2019); Diego (2020); Liao et al. (2020); Mishra et al. (2021); Seo et al. (2022, 2024); Liu et al. (2023); Cheung et al. (2024); Zumalacárregui (2024); Chakraborty and Mukherjee (2024); Chan et al. (2025b), the halo mass function Fairbairn et al. (2023); Tambalo et al. (2023); Çalışkan et al. (2023); Savastano et al. (2023), the structures of different lenses Lo et al. (2025); Vujeva et al. (2025a); Ezquiaga et al. (2025); Vujeva et al. (2025b); Kim et al. (2025); Abe et al. (2025); Ali et al. (2025); Su et al. (2025); Bulashenko et al. (2025); Singh et al. (2025), the cosmic expansion rate Cremonese et al. (2021); Chen et al. (2024); Smith et al. (2025), and the astrophysical environments of GW sources Leong et al. (2025); Li et al. (2025); Ubach et al. (2025), as well as provide an alternate way to test for deviations of general relativity Goyal et al. (2021a, 2023); Narola et al. (2024); Ezquiaga and Zumalacárregui (2020); Liu et al. (2024).
@@
Confirming the lensed nature of a candidate event requires scrutiny, even in small effects such as noise fluctuations and waveform systematics Keitel (2025). Making a confident detection claim requires the simulation and analysis of millions of events. Carefully conducted searches for lensing signatures in GWs in the first three observing runs found no convincing evidence for GW lensing Hannuksela et al. (2019); Dai et al. (2020); Abbott et al. (2021, 2023); Janquart et al. (2023); Chakraborty and Mukherjee (2025). However, the fourth observing run brought a lensed candidate, GW231123 Abac et al. (2025), which we investigate in detail in a companion paper Juno C. L. Chan, J. M. Ezquiaga, R. K. L. Lo, J. Bowman, L. Magaña Zertuche, and L. Vujeva (2025).
@@
Deep-learning techniques have emerged as a powerful solution for performing fast and accurate PE of GW signals Liang and Wang (2025). Unlike common Bayesian sampling methods — which require repeated waveform evaluations and are therefore computationally expensive — neural networks can learn the complex mapping between GW strain data and source parameters directly from simulated training sets. Once trained, these models can generate full posteriors within seconds, achieving orders-of-magnitude speedups while maintaining precision comparable to traditional PE pipelines Chua and Vallisneri (2020); Gabbard et al. (2022); Cuoco et al. (2021); Langendorff et al. (2023); Zhao et al. (2025); Bhardwaj et al. (2023); Cuoco et al. (2025); Hu et al. (2024); Polanska et al. (2024); Stergioulas (2024); Papalini et al. (2025); Santoliquido et al. (2025). Recent studies have begun extending deep-learning techniques to the lensed GW regime, exploring the use of convolutional-, recurrent-, and transformer-based architectures for rapid identification of lensing signatures Kim et al. (2021); Goyal et al. (2021b); Magare et al. (2024) and inferring lens parameters Bada-Nerin et al. (2025); Qin et al. (2025).
@@
Among these approaches, the DINGO framework Green et al. (2020); Green and Gair (2021); Dax et al. (2021a, b); Wildberger et al. (2023); Dax et al. (2023); Gupte et al. (2024); Santoliquido et al. (2025); Hu (2025) represents a state-of-the-art implementation of simulation-based inference using neural posterior estimation (NPE). DINGO combines an embedding network with normalizing flows to produce full Bayesian posteriors conditioned on GW data, enabling rapid and high-fidelity inference across high-dimensional parameter spaces. Simulation-based inference provides a solution by combining computational efficiency and statistical interpretability. Instead of relying on explicit likelihood evaluations, simulation-based inference employs neural density estimators trained on simulated data to learn the complex mapping between observations and parameters Cranmer et al. (2020); Liang and Wang (2025).
@@
Building on top of the simulation-based inference algorithm, DINGO, we present a new lensing inference pipeline DINGO-lensing [^1] https://github.com/dingo-lensing/dingo-lensing to perform NPE. In particular, we employ the geometric optics approximation in the wave optics regime in our lensed waveform generation. By directly learning the posterior distribution from simulated data, DINGO-lensing circumvents the computational bottlenecks of a traditional sampler while maintaining the statistical rigour of conventional inference. This approach establishes a scalable and robust pathway for identifying and characterizing lensed GW events in the era of high-cadence detections.
@@ II DINGO-lensing
In this work, we perform NPE training with the simulation-based inference framework DINGO. We use the fifth observing run (O5) noise amplitude spectral density (ASD) Abbott et al. (2018); sen , assuming Gaussian and stationary noise. Throughout the study, we utilize the inspiral-merger-ringdown phenomenological model with higher modes and precession, IMRPhenomXPHM Pratten et al. (2021) implemented in LALSimulation LIGO Scientific Collaboration et al. (2018); Wette (2020), with a frequency range of $20,1024]\penalty 10000\ \rm{Hz}$ and a waveform duration of $8\penalty 10000\ \rm{s}$. We have modified the network to allow for the use of numerical relativity surrogate waveforms during training and parameter inference, e.g., NRSur7dq4 that spans a frequency range of [20, 512] Hz, which we use in the companion paper Juno C. L. Chan, J. M. Ezquiaga, R. K. L. Lo, J. Bowman, L. Magaña Zertuche, and L. Vujeva (2025). The intrinsic parameters consist of the component masses in the detector frame $m_{i}\penalty 10000\ (i=1,2)$, the spin magnitudes $a_{i}$, and the spin angles $(\phi_{i},\phi_{12},\phi_{JL})$. Other parameters included are the inclination angle $\theta_{JN}$, right ascension (RA) and declination (DEC), polarization angle $\psi$, luminosity distance $d_{\rm L}$, the time of the peak amplitude at geocenter $t_{\rm c}$, and the phase at a reference frequency $\phi_{\rm ref}$ of 20Hz.
@@ II.1 Lensed Waveform Model
With the weak-gravity and thin-lens approximation, the gravitational lensing of a GW signal $h(t)$ can be solved efficiently in the frequency domain Schneider et al. (1992). The lensed $\tilde{h}_{\rm L}(f)$ and unlensed $\tilde{h}_{\rm U}(f)$ frequency-dependent GW signals are related by the amplification factor $F(f)$ as
@@
The amplification factor is obtained by solving the diffraction integral,
$F(f,\bm{\theta}_{\rm S})=\frac{\tau_{D}f}{i}\int\mathrm{d}^{2}\bm{\theta}\,e^{2\pi ift_{\rm d}(\bm{\theta},\bm{\theta}_{\rm S})}\,$, (2)
@@
which integrates over the image plane positions $\vec{\theta}$. Here, $\tau_{D}$ is the characteristic timescale associated to the distances between source-lens-observer, $\bm{\theta}_{\rm S}$ is the nonlensed source angular position, and $t_{\rm d}(\bm{\theta},\bm{\theta}_{\rm S})$ is the time delay between the source and observer, often called the time-delay surface.
@@
Using the stationary phase approximation (SPA), the diffraction integral simplifies to a sum over multiple copies of the original signal. Each of these copies can be magnified, time delayed and phase shifted. If the time delay between any of the repeated copies is less than the signal’s duration, then there are interference effects that distort the signal. In this work we focus on an amplification factor that depends on the time delay $\Delta t$ and relative magnification $\mu_{\rm rel}$ between two images with a phase difference of $\pi/2$, i.e.,
$F(f>0)=1+\sqrt{\mu_{\rm rel}}\,e^{i(2\pi f\Delta t-\pi/2)}\,.$ (3)
@@
This is a good approximation for systems in which two images are interfering with each other and has been shown to hold for several types of singularities, including point, fold, and (some) cusp cases (see Serra and Bulashenko (2025); Ezquiaga et al. (2025) where further details on its validity can be found). Operationally, we implement this lens model coded in modwaveforms. [^2] https://github.com/ezquiaga/modwaveforms
@@
Image properties are determined by their locations on the time delay surface. That is, images that form on minima of the time delay surface are of Type I, those that form on saddle points are of Type II with a $\pi/2$ phase shift, and those on maxima are of Type III with a $\pi$ phase shift. For example, a point-mass lens always produces two images of Type I and II. For the point, fold, and many configurations near cusp singularities, the brightest image (of Type I) arrives before the second brightest image (of Type II). This allows us to define our relative magnification, or the ratio of the second brightest to the brightest image $\mu_{\rm rel}\equiv\mu_{2}/\mu_{1}$, as being $\mu_{\rm rel}\in[0,1]$ and the time delay as $\Delta t>0$. This will always result in a $\pi/2$ phase difference between the two lensed images. This assumption removes a possible degeneracy where images of the same type interfering with each other can lead to $(\mu_{\rm rel},\Delta t)=(0,0)$ or $(\mu_{\rm rel},\Delta t)=(1,0)$, which is broken in our case due to the phase difference between the two images. Therefore, within our framework we define a “lensed image” as an event in which our analysis finds $\Delta t>0$, implying $\mu_{\rm rel}<1$. We define a “nonlensed event” as an event which our analysis finds a relative magnification factor of $\mu_{\rm rel}=0$. Exploring lensing configurations where one would find overlapping images of the same type, as well as images with $\mu_{\rm rel}>1$ such as the ones created by lenses containing substructure Vujeva et al. (2025a, b), is left to future work.
@@ II.2 Training Setup
We adopt the two-stage training scheme with the hyperparameters, where the initial training stage (Stage 0) freezes the initial SVD layer and is unfreezed in the fine-tuning stage (Stage $1$). The training of both the lensed and nonlensed NPE models is performed in two stages. During Stage 0, the model is trained for 300 epochs using the Adam optimizer Kingma and Ba (2017) with a learning rate of $7\times 10^{-5}$ and a cosine learning-rate scheduler Loshchilov and Hutter (2016). In Stage 1, all layers are unfrozen for fine-tuning over an additional 150 epochs, with the learning rate reduced to $1\times 10^{-5}$. Both stages employ a batch size of 4096 and temperatures of 300 and 150, respectively. This two-stage procedure stabilizes early optimization and enables the network to refine the embedding representation for improved posterior accuracy.

### v2

@@ Abstract
Gravitational waves (GWs) can be distorted by intervening mass distributions while propagating, leading to frequency-dependent modulations that imprint a distinct signature on the observed waveforms. Bayesian inference for GW lensing with conventional sampling methods is costly, and the problem is exacerbated by the rapidly growing GW catalog. Moreover, assessing the statistical significance of lensed candidates requires thousands, if not millions, of simulations to estimate the background from noise fluctuations and waveform systematics, which is infeasible with standard samplers. We present a novel method, DINGO-lensing, for performing inference on lensed GWs, extending the neural posterior estimation framework DINGO. By comparing our results with those using conventional samplers, we show that the compute time of parameter estimation of lensed GWs can be reduced from weeks to seconds, while preserving accuracy both in the posterior distributions and the evidence ratios. We train our neural networks with LIGO detector noise at design sensitivity and a lens model that accommodates two overlapping images with a constant $\pi/2$ phase shift. We show that the lensing parameters are recovered with millisecond precision for the time delays. We also demonstrate that our network can identify signals diffracted by point masses, highlighting its flexibility for searches. By simulating thousands of lensed and nonlensed events, we determine how the detectability changes with different source properties. DINGO-lensing provides a scalable and efficient avenue for identifying and characterizing gravitationally lensed GW events in the upcoming observing runs.
@@ I Introduction
GWs propagate over cosmological distances and are not impervious to intervening matter. Instead, their path is affected by nearby gravitational potentials (acting as lenses), such as those of stars, black holes, galaxies, and clusters of galaxies, leading to gravitational lensing Nakamura (1998); Takahashi and Nakamura (2003). In strong gravitational lensing, multiple signals, or images, of the source are produced each with a distinct amplitude, phase shift, and time delay Wang et al. (1996); Dai and Venumadhav (2017); Ezquiaga et al. (2021). The arrival time difference between these images can be very small for smaller lens masses (or specific source-lens-observer configurations) Schneider et al. (1992); Lo et al. (2025); Vujeva et al. (2025a); Ezquiaga et al. (2025); Vujeva et al. (2025b), making it challenging to look for signals coming from the same source. Additionally, these small lens masses may produce overlapping images which causes a beating pattern in the inspiral portion of the waveform resembling precession or the effect of two different binary signals overlapping Liu and Kim (2024); Rao et al. (2025). When the GW wavelength is comparable to or larger than the characteristic scale of the lens, diffraction becomes significant and a single signal can exhibit frequency-dependent modulations. Detecting such signal can be challenging in matched-filtering searches Chan et al. (2025a), but it can answer key questions in the small lens mass regime regarding small compact lens populations Jung and Shin (2019); Lai et al. (2018); Christian et al. (2018); Dai et al. (2018); Diego et al. (2019); Diego (2020); Liao et al. (2020); Mishra et al. (2021); Seo et al. (2022, 2024); Liu et al. (2023); Cheung et al. (2024); Zumalacárregui (2024); Chakraborty and Mukherjee (2024); Chan et al. (2025b), the halo mass function Fairbairn et al. (2023); Tambalo et al. (2023); Çalışkan et al. (2023); Savastano et al. (2023), the structures of different lenses Lo et al. (2025); Vujeva et al. (2025a); Ezquiaga et al. (2025); Vujeva et al. (2025b); Kim et al. (2025); Abe et al. (2025); Ali et al. (2025); Su et al. (2025); Bulashenko et al. (2025); Singh et al. (2025), the cosmic expansion rate Cremonese et al. (2021); Chen et al. (2024); Smith et al. (2025), and the astrophysical environments of GW sources Leong et al. (2025); Li et al. (2025); Ubach et al. (2025), as well as provide an alternate way to test for deviations of general relativity Goyal et al. (2021a, 2023); Narola et al. (2024); Ezquiaga and Zumalacárregui (2020); Liu et al. (2024).
@@
Confirming the lensed nature of a candidate event requires scrutiny, even in small effects such as noise fluctuations and waveform systematics Keitel (2025). Making a confident detection claim requires the simulation and analysis of millions of events. Carefully conducted searches for lensing signatures in GWs in the first three observing runs found no convincing evidence for GW lensing Hannuksela et al. (2019); Dai et al. (2020); Abbott et al. (2021, 2023); Janquart et al. (2023); Chakraborty and Mukherjee (2025). However, the fourth observing run brought a lensed candidate, GW231123 Abac et al. (2025), which we investigate in detail in a companion paper Chan et al. (2025c).
@@
Deep-learning techniques have emerged as a powerful solution for performing fast and accurate PE of GW signals Liang and Wang (2025). Unlike common Bayesian sampling methods—which require repeated waveform evaluations and are therefore computationally expensive—neural networks can learn the complex mapping between GW strain data and source parameters directly from simulated training sets. Once trained, these models can generate full posteriors within seconds, achieving orders-of-magnitude speedups while maintaining precision comparable to traditional PE pipelines Chua and Vallisneri (2020); Gabbard et al. (2022); Cuoco et al. (2021); Langendorff et al. (2023); Zhao et al. (2025); Bhardwaj et al. (2023); Cuoco et al. (2025); Hu et al. (2024); Polanska et al. (2024); Stergioulas (2024); Papalini et al. (2025); Santoliquido et al. (2025).
Recent studies have begun extending deep-learning techniques for the rapid identification of strongly lensed GW signals Goyal et al. (2021b); Magare et al. (2024). In the wave-optics regime, several architectures have been explored, including visual geometry group networks Kim et al. (2021), conditional variational autoencoders Bada-Nerin et al. (2025), and neural spline flows Qin et al. (2025). Reference Kim et al. (2021) presents a proof-of-principle demonstration that deep-learning models can identify waveform distortions. Reference Bada-Nerin et al. (2025) further shows a substantial speed-up in inferring lens parameters compared to traditional Bayesian sampling. Reference Qin et al. (2025) extends the inference dimensionality to eleven. Among these approaches, however, a rapid inference of all source parameters with accuracy comparable to traditional Bayesian sampling methods has not yet been achieved.
@@
The DINGO framework Green et al. (2020); Green and Gair (2021); Dax et al. (2021a, b); Wildberger et al. (2023); Dax et al. (2023); Gupte et al. (2024); Santoliquido et al. (2025); Hu (2025) represents a state-of-the-art implementation of simulation-based inference using neural posterior estimation (NPE). DINGO combines an embedding network with normalizing flows to produce full Bayesian posteriors conditioned on GW data, enabling rapid and high-fidelity inference across high-dimensional parameter spaces Green and Gair (2021); Dax et al. (2021a). Simulation-based inference provides a solution by combining computational efficiency and statistical interpretability. Instead of relying on explicit likelihood evaluations, simulation-based inference employs neural density estimators trained on simulated data to learn the complex mapping between observations and parameters Cranmer et al. (2020); Liang and Wang (2025). This offers a promising deep-learning pathway toward rapid and accurate inference on lensed GWs.
@@
Building on top of the simulation-based inference algorithm, DINGO, we present a new lensing inference pipeline DINGO-lensing [^1] https://github.com/dingo-lensing/dingo-lensing. In particular, we employ a flexible two image model in our lensed waveform generation that is able to accommodate the most common wave-optics models Ezquiaga et al. (2025). By directly learning the posterior distribution from simulated data, DINGO-lensing circumvents the computational bottlenecks of a traditional sampler while maintaining the statistical rigour of conventional inference. This approach establishes a scalable and robust pathway for identifying and characterizing lensed GW events in the era of high-cadence detections.
@@ II DINGO-lensing
In this work, we perform NPE training with the simulation-based inference framework DINGO. We use the fifth observing run (O5) noise amplitude spectral density (ASD) Abbott et al. (2018); sen , assuming Gaussian and stationary noise. Since glitches cause the noise to deviate from the Gaussian and stationary assumptions used here, we will assess their influence in future studies. Throughout the study, we utilize the inspiral-merger-ringdown phenomenological model with higher modes and precession, IMRPhenomXPHM Pratten et al. (2021) implemented in LALSimulation LIGO Scientific Collaboration et al. (2018); Wette (2020), with a frequency range of $20,1024]\penalty 10000\ \rm{Hz}$ and a waveform duration of $8\penalty 10000\ \rm{s}$. We have modified the network to allow for the use of numerical relativity surrogate waveforms during training and parameter inference, e.g., NRSur7dq4 that spans a frequency range of [20, 512] Hz, which we use in the companion paper Chan et al. (2025c). The intrinsic parameters consist of the component masses in the detector frame $m_{i}\penalty 10000\ (i=1,2)$, the spin magnitudes $a_{i}$, and the spin angles $(\phi_{i},\phi_{12},\phi_{JL})$. Other parameters included are the inclination angle $\theta_{JN}$, right ascension (RA) and declination (DEC), polarization angle $\psi$, luminosity distance $d_{\rm L}$, the time of the peak amplitude at geocenter $t_{\rm c}$, and the phase at a reference frequency $\phi_{\rm ref}$ of 20Hz.
@@ II.1 Lensed waveform model
With the weak-gravity and thin-lens approximation, the gravitational lensing of a GW signal $h(t)$ can be solved efficiently in the frequency domain Nakamura (1998); Takahashi and Nakamura (2003). The lensed $\tilde{h}_{\rm L}(f)$ and unlensed $\tilde{h}_{\rm U}(f)$ frequency-dependent GW signals are related by the amplification factor $F(f)$ as
@@
The amplification factor is obtained by solving the diffraction integral Takahashi and Nakamura (2003),
$F(f,\bm{\theta}_{\rm S})=\frac{\tau_{D}f}{i}\int\mathrm{d}^{2}\bm{\theta}_{\rm{I}}\,e^{2\pi ift_{\rm d}(\bm{\theta}_{\rm{I}},\bm{\theta}_{\rm S})}\,$, (2)
@@
which integrates over the image plane positions $\bm{\theta}_{\rm{I}}$. Here, $\tau_{D}$ is the characteristic timescale associated to the distances between source-lens-observer, $\bm{\theta}_{\rm S}$ is the nonlensed source angular position, and $t_{\rm d}(\bm{\theta}_{\rm{I}},\bm{\theta}_{\rm S})$ is the time delay between the source and observer, often called the time-delay surface. The notations follow the definition in Ezquiaga et al. (2025).
@@
Using the stationary phase approximation (SPA), the diffraction integral simplifies to a sum over multiple copies of the original signal. Each of these copies can be magnified, time delayed and phase shifted. If the time delay between any of the repeated copies is less than the signal’s duration, then there are interference effects that distort the signal. In this work, we focus on an amplification factor that depends on the time delay $\Delta t$ and relative magnification $\mu_{\text{rel}}$ between two images with a phase difference of $\pi/2$, i.e.,
$F(f>0)=1+\sqrt{\mu_{\text{rel}}}\,e^{i(2\pi f\Delta t-\pi/2)}\,.$ (3)
@@
This expression, which is equivalent to the strong-lensing regime Dai and Venumadhav (2017); Ezquiaga et al. (2021), is a good approximation for systems in which two images are interfering with each other and has been shown to hold for several types of singularities, including point, fold, and (some) cusp cases (see Serra and Bulashenko (2025); Ezquiaga et al. (2025) where further details on its validity can be found). Operationally, we implement this lens model coded in modwaveforms. [^2] https://github.com/ezquiaga/modwaveforms
@@
Image properties are determined by their locations on the time delay surface. That is, images that form on minima of the time delay surface are of type I with no phase shift, those that form on saddle points are of type II with a $\pi/2$ phase shift, and those on maxima are of type III with a $\pi$ phase shift Blandford and Narayan (1986). For example, a point-mass lens always produces two images of type I and II Schneider et al. (1992). For the point, fold, and many configurations near cusp singularities, the brightest image (of type I) arrives before the second brightest image (of type II) Blandford and Narayan (1986). Because of this hierarchy in the magnifications, this allows us to define our relative magnification, or the ratio of the second brightest to the brightest image $\mu_{\text{rel}}\equiv\mu_{2}/\mu_{1}$, as being restricted to $\mu_{\text{rel}}\in[0,1]$ and the time delay between the second and first image as $\Delta t>0$. Because the type I image arrives before the type II image, this will always result in a $\pi/2$ phase difference between the two lensed images. This assumption removes a possible degeneracy where images of the same type interfering with each other can lead to $(\mu_{\text{rel}},\Delta t)=(0,0)$ or $(\mu_{\text{rel}},\Delta t)=(1,0)$, which is broken in our case due to the phase difference between the two images. Therefore, within our framework we define a “lensed image” as an event in which our analysis finds $\Delta t>0$, implying $\mu_{\text{rel}}<1$. We define a “nonlensed event” as an event which our analysis finds a relative magnification factor of $\mu_{\text{rel}}=0$. Exploring lensing configurations where one would find overlapping images of the same type, as well as images with $\mu_{\text{rel}}>1$ such as the ones created by lenses containing substructure Vujeva et al. (2025a, b), is left to future work.
@@ II.2 Training setup
Training setting Stage 0 Stage 1
Layers SVD layer frozen All layers unfrozen
Epochs 300 150
Optimizer Adam Kingma and Ba (2017) Adam
Learning rate $7\times 10^{-5}$ $1\times 10^{-5}$
Scheduler Cosine scheduler Loshchilov and Hutter (2016) Cosine scheduler
Batch size 4096 4096
Temperature 300 150
Purpose Stabilize early training Fine-tuning for accuracy
Table 2: Summary of the two-stage NPE training scheme.
We adopt the two-stage training scheme with the hyperparameters summarized in Table 2. The initial training stage (Stage 0) freezes the initial SVD layer and is unfreezed in the fine-tuning stage (Stage $1$). The training of both the lensed and nonlensed NPE models is performed in two stages. During Stage 0, the model is trained for 300 epochs using the Adam optimizer Kingma and Ba (2017) with a learning rate of $7\times 10^{-5}$ and a cosine learning-rate scheduler Loshchilov and Hutter (2016). In Stage 1, all layers are unfrozen for fine-tuning over an additional 150 epochs, with the learning rate reduced to $1\times 10^{-5}$. Both stages employ a batch size of 4096 and temperatures of 300 and 150, respectively. This two-stage procedure stabilizes early optimization and enables the network to refine the embedding representation for improved posterior accuracy.
@@
Note that we tested training configurations with different batch sizes, temperatures, and learning rates. While our network validation in Section III focuses on the setup presented in Table 2, we find that the choice of batch size and temperature introduces negligible variation in performance. A smaller learning rate, however, provides more stable training and better performance when the range of luminosity distance is increased.